OpenStudy (anonymous):

5x+3y=-11 7x-2y=19 please solve using elimination method

OpenStudy (anonymous):

multiply first by 2 and second by 3 then add the resulting equations to eliminate term with x you will get value of y put it in your original first equation and get value of x

OpenStudy (anonymous):

if u need more assistance, let me know....

OpenStudy (angela210793):

multiply the 1st with a 2--->10x+6y=-22 multiply the 2nd with a 3--->21x-6y=57 now add both parts and u'll have 31x=35 x=35/31 then replace x in the 1st or 2nd eq....5*35/31 +3y=-11..and find y

OpenStudy (anonymous):

yes angela has done the first part for u........

OpenStudy (anonymous):

ok so I don't understand how to solve for y?

OpenStudy (anonymous):

and I don't understand how afraction became the answer

OpenStudy (anonymous):

u got x = 35/31 put it in 5x + 3y = -11 5*35/31 + 3y = -11 175/31 + 3y = -11 3y = -11 - 175/31 3y = (-341-175)/31 3y = -516/31 y = -516/93

OpenStudy (angela210793):

x=35/31..then we replace the value of x we found at one of the eq....I chose 1st... 5*35/31 +3y=-11 3y=-11-175/31 y=(-341-175)/93=-516/93

OpenStudy (anonymous):

so you don't have to multiply both sides by the first answer?

OpenStudy (angela210793):

noo

OpenStudy (anonymous):

I just always get confused on what steps to take for what problems. - numbers throw me off too. Thank you both for your help.

OpenStudy (anonymous):

MY assignment say those answers are wrong....now I am even more confused because it started making a little bit of sense

OpenStudy (angela210793):

I don't see any mistake in the solution....

OpenStudy (anonymous):

the answer should look like this right. (35/31,-156/93) ?

OpenStudy (angela210793):

yea but's -516 not -156

OpenStudy (anonymous):

ok that's my fault...hold on.

OpenStudy (anonymous):

yea still says wrong???

OpenStudy (anonymous):

it might be a typo error in the printing...

OpenStudy (anonymous):

should the first answer be be a -?

OpenStudy (anonymous):

It says the second part should be -172/31

OpenStudy (anonymous):

so it should look like this (35/31,-172/31)

OpenStudy (anonymous):

yes this is right, we forgot to reduce the value of y to lowest terms -516/93 = -172/31 (divide both numerator and denominator by 3)

OpenStudy (anonymous):

now do it and check, our working is right...

OpenStudy (anonymous):

now u can push THE GOOD ANSWER button.... ( :))

OpenStudy (anonymous):

??????????

OpenStudy (anonymous):

LOL I always give you a good answer!

OpenStudy (anonymous):

thanks !!!

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

you help me I give you medals every single time!

OpenStudy (anonymous):

i am lucky to b appreciated

OpenStudy (anonymous):

I am lucky for the help. Angela I gave you a medal as well.

OpenStudy (angela210793):

Lucky u..I solve the q and ppl say thnx Harkirat....

OpenStudy (angela210793):

yea I saw tht..( was having dinner ) thnx

OpenStudy (anonymous):

I thank you as well both of you helped me very much. I thought you were gone.

OpenStudy (angela210793):

Nnoononono...Wht I said wasn't 4 u Aroub :):):):)

OpenStudy (anonymous):

hey angela don't get jealous :D

OpenStudy (angela210793):

I'm not jealous....and I don't really care abt the medals...I just want 4 ppl to be polite and say a **thanks**

OpenStudy (anonymous):

generally they thank all who hv contributed, and so has plshlpme !!

OpenStudy (anonymous):

I was going to thank you, I was just responding to Harkirat. I apologize. I didn't mean to make you fell unappreciated because you are.

OpenStudy (angela210793):

Errr......I said I wasn't talking abt Aroub.....

OpenStudy (angela210793):

Aroub HELLO!!!!!! the words I said weren't 4 u

OpenStudy (anonymous):

ok guys. I appreciate you both again THANK YOU BOTH FOR YOUR HELP! MUCH APPRECIATED